OpenStudy (sarahc):

Find a ⋅ b. a = 6i - 6j, b = 4i + 2j 12 -36 <24, -12> <10, -4>

OpenStudy (swissgirl):

Multiply the i components and mutliply the j components and then add together So its (6*4)+(-6*2)=24-12=12

OpenStudy (swissgirl):

all i see are a bunch of triangles

OpenStudy (sarahc):

Blah, me too… The given measurements may or may not determine a triangle. If not, then state that no triangle is formed. If a triangle is formed, then use the Law of Sines to solve the triangle, if it is possible, or state that the Law of Sines cannot be used. B = 111°, c = 8, b = 12 A. C = 38.5°, A = 30.5°, a ≈ 6.5 B. No triangle is formed. C. C = 30.5°, A = 38.5°, a ≈ 6.5 D. The triangle cannot be solved with the Law of Sines.

OpenStudy (sarahc):

Idk why it does that..

OpenStudy (sarahc):

the triangles aren't supposed to be there for example… C = 30.5, A = 38.5, a 6.5

OpenStudy (swissgirl):

ok now i see it

OpenStudy (sarahc):

same…..weird

OpenStudy (swissgirl):

Its annoying cuz I dont have a scientific calc on hand ....

OpenStudy (sarahc):

oh..:/ thats alright. thanks for your help